A catering company needs to provide 10 lbs of mixed nuts for a wedding reception. The couple has budgeted $42 for nuts. Peanuts cost $3.00/ lb and cashews cost $7.00/ lb. how many pounds each of peanuts and cashews should the catering company combine?
Answer by checkley79(3341)

7x+3(10-x)=42
7x+30-3x=42
4x=42-30
4x=12
x=12/4
x=3 pounds of cashews are used.
10-3=7 pounds of peanut are used.
proof:
7*3+3*7=42
21+21=42
42=42